A box contains 6 white balls and 6 black balls. Find number of ways 4 balls can be
drawn from the box if
i) Two must be white
ii) All of them must have same color
1
Expert's answer
2020-11-02T19:57:49-0500

i) if we choose 2 white balls, then we choose 2 black balls too.

The number of ways of choosing 2 white(also for black) balls from 6 is: "C^{2}_6"

Then, the number of ways of choosing 2 white and 2 black balls is:

"C^{2}_6*C^{2}_6=\frac{6!}{2!*4!}*\frac{6!}{2!*4!}=225"

ii) All chosen balls can be either black or white.

The number of ways of choosing 4 white(also for black) balls from 6 is: "C^{4}_6"

Then, the number of ways of choosing 4 white or 4 black balls is:

"C^{4}_6+C^{4}_6=2*C^{4}_6=2*\frac{6!}{2!*4!}=30"